OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 + 8x − 1. (−2, 7) (2, 7) (2, −9) (−2, −9)

OpenStudy (anonymous):

hmm same as before \[-\frac{b}{2a}=-\frac{8}{2\times (-2)}=2\]

OpenStudy (anonymous):

second coordinate is what you get when you replace \(x\) by \(2\)

OpenStudy (anonymous):

kkk :)

OpenStudy (anonymous):

the last one :)

OpenStudy (anonymous):

if you need help with that part let me know but it should be straight forward

OpenStudy (anonymous):

i mean the third one :)

OpenStudy (anonymous):

ugh the second one :)

OpenStudy (anonymous):

careful

OpenStudy (anonymous):

yes, the second one

OpenStudy (anonymous):

thanks .. please continue helping me :)